### v1

# Entanglement signatures of a percolating quantum system

Abstract
Entanglement measures have emerged as one of the versatile probes to diagnose quantum phases and their transitions. Universal features in them expand their applicability to a range of systems, including those with quenched disorder. In this work, we show that when the underlying lattice has percolation disorder, free fermions at a finite density show interesting entanglement properties. These emerge from massively degenerate eigenspaces where the notion of standard bipartite entanglement entropy needs to be revisited. We show that appropriately defined entanglement measures can capture both geometrical aspects and electronic correlations of the percolated quantum system. Using a variety of toy models, we uncover the rich phenomenology of this physics. We further show that there are decipherable signatures of the classical percolation transition in these entanglement properties. Our work presents an interesting interplay between the geometrical properties of a lattice and quantum entanglement in a many-body quantum system.
Introduction: Quantum entanglement [1, 2] and their measures [3, 4, 5] have established themselves as necessary tools to diagnose quantum phases and their transitions in a variety of systems, be it clean or disordered[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The universal features in quantities such as bipartite entanglement entropy [3] and entanglement spectrum [17] have also made it a defining quantity for studying unconventional quantum phases such as topological phases [18, 19, 17, 20, 21, 22, 23], particularly in presence of disorder [24, 25]. Among disordered systems, percolation problems [26, 27, 28, 29], where a lattice is probabilistically diluted either on the sites or bonds, are known to exhibit second-order phase transitions, namely geometrical phase transition, with universal critical exponents [30, 31, 32, 29, 33]. The critical point is known as the classical percolation threshold $p_{c}$, where the lattice suddenly gets geometrically disconnected [29]. The interplay of such geometric transitions has been long studied in both classical and quantum systems [34, 35, 36, 37, 38, 39] and recently in the context of topological phases [40, 41, 42]. However, the interplay of percolation disorder and entanglement properties has been little explored [43]. We visit entanglement measures in light of the percolation disorder in the simplest of fermionic quantum systems: free fermions hopping on a lattice. In particular, we address do entanglement measures show signatures of a geometrical phase transition? Do they still follow the usual area/volume law [44, 45] diagnostics?
In this work, we investigate above questions to show that in percolating quantum systems, the conventional notions of bipartite entanglement properties do not hold. Since we will focus on the free fermionic systems, it is expected that entanglement entropy should be straightforwardly calculated using the Peschel prescription of correlator matrices [46, 47]. However, as we will show in percolation disorder, such calculations have quite a few subtleties. The non-intuitive aspect of the results, particularly in entanglement, arises from massive exact degeneracy due to lattice percolation. In particular, we show, that one needs to investigate three distinct quantities: typical entanglement ($S_{\text{typ}}$), annealed entanglement ($S_{\text{ann}}$) and quenched entanglement ($S_{\text{quen}}$) each of which captures distinctive signatures. Interestingly, they often depend on the number of exact zero-modes ($N_{0}$) in the system, which is in turn related to the geometrical aspects of the lattice. We finally show that even the classical percolation threshold has distinctive signatures, where entanglement scaling depends on the emergent fractal nature of the largest cluster.
Measures of entanglement and free fermions: Given a wavefunction $|\psi\rangle$, or the density matrix ($\hat{\rho}=|\psi\rangle\langle\psi|$) defined for a system composed of two subsystems $A$ and $B$, the entanglement entropy of region $A$ with $B$ (or vice-versa) is given by $S_{A}=-\text{Tr}\big{(}{\hat{\rho}_{A}\log\hat{\rho}_{A}}\big{)}$, where $\hat{\rho}_{A}=\text{Tr}_{B}\hat{\rho}$. However, this definition assumes that the system is described by a unique wavefunction $|\psi\rangle$. But in presence of degeneracies, it is more appropriate to study typical entanglement
$S_{\text{typ}}=\langle S_{A}(\hat{\rho}_{A})\rangle,$ (1)
where the averaging is done over pure state ensembles made of degenerate wavefunctions [48, 49]. The wavefunction coefficients are complex and drawn from normal probability distributions implementing a uniform Haar measure over the unitary transformations about any reference state [50, 51, 52, 53]. $S_{\text{typ}}$ is however different from
$S_{\text{ann}}=S_{A}(\langle\hat{\rho}_{A}\rangle),$ (2)
where the averaging on the reduced density matrix is done before evaluating its entanglement content. Such a measure we call annealed entanglement entropy, where the annealed average renders $\langle\hat{\rho}_{A}\rangle$ to be of a mixed character. In general $S_{\text{typ}}<S_{\text{ann}}$, since the latter reflects the maximum information content possible in a given Hilbert space consistent with Page’s result [48]. For instance, in a decoupled two spin system while $S_{A}$ may seem to be zero, $S_{\text{typ}}\sim 0.33$, $S_{\text{ann}}=\log(2)$ (for details see Supplemental Material (SM) [54]).
In a many-body free fermionic system defined on $N$ sites such that $c^{\dagger}_{i}$ and $c_{i}$ are the fermionic creation and annihilation operators, the ground state is often a unique Slater determinant and the bipartite entanglement is evaluated using the Peschel formulation of correlator matrix ${\cal C}$ where its elements $c_{ij}=\langle c^{\dagger}_{i}c_{j}\rangle$[46, 47]. However, for a system with ground state degeneracies, the measure of typical and annealed entanglement becomes important. Thus while for any single Slater determinant, the entanglement entropy can be given by the corresponding correlator matrix $S({\cal C}_{A})$, when averaged over all choices of such random coefficients over the degenerate manifold, one gets $S_{\text{typ}}=\langle S({\cal C}_{A})\rangle$ and similarly $S_{\text{ann}}=S(\langle{\cal C}_{A}\rangle)$ (for an illustrative example of these quantities in a four-site toy model, see SM [54]). A simple way to break such exact degeneracies would be to add a small perturbative Anderson disorder $\eta\sim 10^{-12}$, which will choose a unique ground state. Then we define quenched entanglement, to capture the fermionic correlations in the system
$S_{\text{quen}}=\langle S({\cal C}_{A})\rangle_{\eta},$ (3)
where the average is over disordered configurations. Before delving into percolation problems, we illustrate the role of these measures in a zero-dimensional system.
Figure 1: Zero dimensional model: (a) $S_{\text{quen}}$ (averaged over $N_{C}\equiv 10^{3}$ configurations) grows linearly with number of sites, $N$. (inset) Schematic of the model and partitions $A$ and $B$. (b) Behaviour of $S_{\text{typ}}$ and $S_{\text{ann}}$ with $N$. The number of zero-modes, $N_{0}=N-1$. Both are calculated using $N_{R}\equiv 10^{4}$ ensembles of random pure states.
Zero-dimensional system: Consider a $N$ site Hamiltonian where the fermions can hop from a site to any other site with strength $t=-1$ (see Fig. 1(a) inset) such that $H=-\sum_{i,j}c^{\dagger}_{i}c_{j}$. The single particle spectrum has one eigenvalue with energy $-N$, while $N-1$ eigenvalues are exactly $zero$. A half-filled system here is thus ${}^{N-1}C_{\frac{N}{2}-1}$ fold degenerate. An infinitesimal $\eta$ disorder of strength $\eta=10^{-12}$ can be added in form of a term $\sum_{i}\epsilon_{i}c^{\dagger}_{i}c_{i}$ where $\epsilon\in[-\eta,\eta]$ to break this exact degeneracy. Given the long-range character of hopping, the entanglement content for equal bipartition of the system such that $N_{A}=N/2$ is still volume law where $S_{\text{quen}}=s_{v}N+s_{0}$ with a coefficient $s_{v}\sim 0.002$ as shown in Fig. 1(a). However, when such massive degeneracy is intact - that reflects in
$S_{\text{typ}}=S_{\text{quen}}+(\log 2-\frac{1}{2})(N_{0}-1),$ (4)
where the second term arises from the effective geometrical component because of $N_{0}=N-1$ number of zero-modes of the system (see Fig. 1(d)). The factor of $(\log 2-\frac{1}{2})$ arises from the effective random pure states made out of $N_{0}$ zero-modes [55]. In general, when the number of occupied zero-modes is $fN_{0}$ ($10<f<$ 1), the geometrical volume law is empirically proportional to $f(1-f)$ representing the effective phase space volume (see SM [54]). Another alternate measure of the entanglement in this system is to average the correlator matrix first, which in general, corresponds to a mixed-density matrix. For such a system while $c_{ii}=\frac{1}{2}$ and $c_{ij}=\frac{1}{2N}$ [56] leading to $S_{\text{ann}}\sim\frac{N_{0}}{2}\log 2.$
While it may appear that all entanglement measures are volume law and therefore similar, it is pertinent to emphasize that all three quantities represent three distinct entanglement measures having different physical content. While $S_{\text{quen}}$ captures the fermionic correlations, which are inherently volume law since the network is zero-dimensional, the character of $S_{\text{typ}}$ and $S_{\text{ann}}$ capture the massive degeneracy of the system - where $S_{\text{typ}}$ measures the average entanglement for any choice of a typical pure state, $S_{\text{ann}}$ measures the entanglement due to a mixed state arising from averaging the density matrix first. For instance, the entanglement measure of the full system has $S_{\text{ann}}\neq 0$ but $S_{\text{typ}}=0$. Having discussed the subtleties and the different measures of entanglement, we now discuss quantum percolation problems, where studying these various measures of entanglement become indispensable given spectral degeneracies.
One-dimensional percolation: We first discuss percolation in one-dimensional lattice, where spinless fermions hop with the following tight-binding Hamiltonian
[…]
where $L$ is the system size and $t=1$. The probability of having a bond is given by $p$ such that at $p=0$, the system contains a completely decoupled set of sites, while at $p=1$ it is a translationally invariant one-dimensional chain (see Fig. 2(a)). The percolation transition, where the connectivity of one end of the lattice gets disconnected from the other end, happens at $p=p_{c}=1$ [31, 57]. The fermion filling is kept fixed at $=1/2$. Given the Hamiltonian is real and has a sublattice symmetry, it belongs to the BDI symmetry class [58, 59], which is retained under the percolation protocol.
The fermionic ground state describes a Fermi sea at $p=1$; however, at any $p<1$, given the fermions reside on disconnected clusters, the ground state should be interpreted as an Anderson insulator state. This is consistent with the effect of any uncorrelated disorder in one dimension[60, 61]. The entanglement content, therefore, is generically expected to be $\sim\log L$ at $p=1$ and $\mathcal{O}(1)$ in presence of disorder, as is known from Cardy-Calabrese result for critical states [10, 62, 12] and area law entanglement for short-range correlated states [44, 14]. At any $p$ given the presence of disconnected clusters, there are spectral degeneracies that can be split using an infinitesimal disorder $\eta$ to obtain $S_{\text{quen}}$ (see Fig. 2(b)). An analytical estimate can be obtained as follows. Given $P_{s}$ is the probability of having a $s$ cluster, in general, its maximal entanglement content in equal bi-partition is ($\frac{c}{6}\log(\frac{s}{\pi})+c_{0}$) [10] where $c$ is the central charge ($c=1$) of the one-dimensional bosonic CFT and $c_{0}$ is an area law piece. Thus for a thermodynamic system,
$\tilde{S}_{\text{quen}}=\sum_{s=2}^{\infty}P_{s}\Big{(}\frac{c}{6}\log(\frac{s}{\pi})+c_{0}\Big{)}$ (6)
where $\tilde{S}_{\text{quen}}$ represents a configuration averaged value over $S_{\text{quen}}$. Here $P_{s}=sn_{s}=s(1-p)^{2}p^{s-1}$ which, given any $p$, is the probability that an arbitrary site of the diluted chain belongs to a cluster of $s$ number of sites [29]. This has a linear rise at small $p$ and a divergence near $p=1$. This analytic behaviour, along with our numerical results, is shown in Fig. 2(b)(here, $c_{0}=0.409\pm 0.002$). Broadly the behaviour indeed remains area law (with a $p$ dependent coefficient) except at $p=1$ where the logarithmic $L$ dependence is restored. Interestingly the average cluster size $\langle M\rangle=\sum_{s}sP_{s}$, diverges near $p\rightarrow p_{c}$ as $(p_{c}-p)^{-\gamma}$ with $\gamma=1$. Thus $\tilde{S}_{\text{quen}}\sim\frac{c}{6}\log\langle M\rangle\sim\frac{c}{6}\log\xi^{\frac{\gamma}{\nu}}$, where the geometric correlation length $\xi$ also diverges as $(p_{c}-p)^{-\nu}$ with $\nu=1$. Since at $p=1$, $\xi\sim L$, one expects a scaling where, $e^{6\tilde{S}_{\text{quen}}}L^{-\frac{\gamma}{\nu}}\sim 1$ which is shown in Fig. 2(c). Thus, the entanglement measure captures the percolation exponents near the geometrical phase transition here at $p_{c}=1$. However, these results, as discussed before, required us to put an infinitesimal degeneracy splitting disorder $\eta$, which also breaks the symmetry of the full Hamiltonian. In absence of any such disorder, the massive degenerate manifold of zero-modes leads to geometrical components of entanglement entropy, as we discuss next.
As the one-dimensional lattice is percolated, various clusters of different sizes appear on the chain. For any odd $s$ the cluster has one single particle zero energy mode. Thus, the zero-mode density is,
$\frac{{N}_{0}}{L}=\sum_{m=0}^{\infty}n_{2m+1}=\frac{1-p}{1+p},$ (7)
for a $L$ sized chain at percolation probability $p$. This analytical behaviour and the disorder averaged numerical estimate of zero-modes number $\tilde{N_{0}}$ are shown in Fig. 2(e). Now, a half-filled state in such a system again leads to highly degenerate many-body eigenspace. A uniform Haar measure here leads to
$\tilde{S}_{\text{typ}}=\tilde{S}_{\text{quen}}+(\log 2-\frac{1}{2})\tilde{N}_{0},$ (8)
while $\tilde{S}_{\text{ann}}\sim\frac{\tilde{N}_{0}}{2}\log 2$. All the behaviours match the numerical results as shown in Fig. 2(d). Thus, geometrical disorder, as in one-dimensional percolation, provides distinctive signatures in various entanglement measures both from intra-cluster fermionic correlations and from geometric components of the lattice itself.
Figure 2: Bond diluted chain: (a) Subsystems $A$ and $B$ in a $L$ sized chain, with bond occupation probability $p$. (b) $\tilde{S}_{\text{quen}}$ with $p$ for different values of $L$ (c) Scaling of $\tilde{S}_{\text{quen}}$ near $p=p_{c}=1$ with $\gamma=\nu=1$. (d) Behaviuor of $\tilde{S}_{\text{typ}}$ and $\tilde{S}_{\text{ann}}$ with $p$ (e) Zero-mode density ($\equiv\tilde{N}_{o}/L$) for $L=600$ compared to analytical result [63]. In (b), (c), and (e) $N_{C}=10^{3}$, in (d) $N_{C}=40$ and for each configuration, $N_{R}=10^{2}$.
Figure 3: Bond diluted square lattice: (a) $\tilde{S}_{\text{quen}}$ with $p$ for different system sizes. ($N_{C}=10^{2}$). (inset) Schematic of a configuration with A and B partitions. (b) $\tilde{S}_{\text{typ}}$ and $\tilde{S}_{\text{ann}}$ with $p$. $N_{C}=10,N_{R}=10^{2}$ (c) $\tilde{N}_{0}$ with $p$ ($N_{C}=10^{2}$), compared to lattice animal results (see text). In (b-c), $L=48$.
Two-dimensional percolation: In two dimensional percolation the system has a finite $p_{c}$; for instance in square lattice bond percolation it is known $p_{c}=\frac{1}{2}$ [64, 65]. At $p=p_{c}$ a spanning cluster develops with critical exponent $\gamma=43/18$ [29]. We again pose the question of different entanglement measures for this system. As is known that for a two-dimensional free fermionic system - any infinitesimal disorder localizes all the wavefunctions [60, 61], thus in terms of electronic properties, we expect the system to be localized for all values of $p$ [66, 67, 68, 69] even though there have been studies finding numerical evidence otherwise [70, 71, 72, 73, 74, 75, 76]. Given any $p$ the mean number of clusters containing $k$ bonds, $n_{k}$ is given by $n_{k}=\sum_{t}g(k,t)p^{k}(1-p)^{t}$ where $t$ is the perimeter of the cluster and $g(k,t)$ is the geometrical factor associated with the number of lattice animals given ($k,t$) [77]. Since any square lattice is a bipartite graph with a symmetric spectrum, the $E_{F}$ remains pinned to zero even under percolation. We find that $S_{\text{quen}}$ follows an area law behaviuor (see Fig. 3(a)) i.e. $\propto L$ for $p<1$. While at $p=1$, the complete square lattice is restored, leading to a finite Fermi sea, the entanglement is proportional to $L\log L$ given by the Widom conjecture [12, 78]. At $p<1$, however, such a momentum space description is no longer applicable. At small $p$, one can enumerate the lattice animals exactly and count their entanglement contribution, as shown by the analytical curve in Fig. 3(a) (for details, see SM [54]). This is in contrast to $S_{\text{typ}}$ and $S_{\text{ann}}$, which again depend on the extensive number of zero-modes present in the system (see Fig. 3(b)). The density of zero-modes can again be estimated analytically from lattice animals, given by
$\frac{N_{0}}{L^{2}}=\sum_{k,t}n_{0}(k,t)g(k,t)p^{k}(1-p)^{t},$ (9)
where $n_{0}(k,t)$ is the number of zero-modes in a cluster of bond-size $k$ and perimeter $t$. A lower bound on $N_{0}$, calculated using lattice animals up to $k=4$, is shown in Fig. 3(c), and it matches well with disorder averaged zero-mode number $\tilde{N}_{0}$ for small $p$ values (details in SM [54]). Given the zero-modes really appear from the geometrical aspects of the clusters, it is thus imperative that they also determine the entanglement content.
Figure 4: Largest cluster: (a) Disorder averaged $S^{lc}_{\text{quen}}$ with $p$, for square lattice of different size $L$ ($N_{c}=10^{2}$). (insets) typical configurations for two values of $p$ where dark-blue sites form the largest cluster. (b) Scaling of $S^{lc}_{\text{quen}}$ with the fractal dimension $d_{f}=91/48$ shows crossing at percolation threshold $p_{c}=0.5$. In (c-d) similar to (a-b) but for triangular lattice with $p_{c}\sim 0.35$ (see text).
To distil any signature of percolation transition at $p=p_{c}$, we calculate the configuration averaged (quenched) bipartite entanglement entropy of the largest cluster ($\equiv S^{lc}_{\text{quen}}$) as a function of $p$ and show this in Fig. 4(a). Interestingly, while for $p<p_{c}$ $S^{lc}_{\text{quen}}\propto L^{0}$, for $p>p_{c}$ $S^{lc}_{\text{quen}}\propto L$ where $L$ is the linear dimension of the lattice. Given it is known that the largest cluster follows a scaling $L^{d_{f}}$ [29] where $d_{f}$ is the fractal dimension of the system, for a typical area law behaviour, we expect $S^{lc}_{\text{quen}}\propto L^{d_{f}/2}$ near $p_{c}$. A scaling collapse using this form shows a crossing (see Fig. 4(b)) at $p=p_{c}=0.5$ with $d_{f}=91/48$ as known for two-dimensional percolation transition. Since this physics should be independent of the microscopic lattice, we apply the same analysis to a tight-binding triangular lattice (Fig. 4(c)). Again, the same collapse shows a crossing (Fig. 4(d)) near $p=p_{c}=2\sin(\frac{\pi}{18})$ [64] illustrating that the entanglement scaling indeed follows the universal features of geometrical phase transitions even though the exact value of $p_{c}$ is itself not universal.
Outlook: Quantum entanglement and its measures have taken a defining role in deciphering the nature of quantum phases. In particular, the nature of low energy excitations are often equated with whether the bipartite entanglement entropy follows area law, or has logarithmic corrections. In this work, we revisit various measures of quantum entanglement in the context of percolation disorder in free fermionic lattice Hamiltonians. We find that percolation disorder inherently generates extensive degeneracies, which gives rise to subtleties in standard bipartite entanglement entropy. It is then important to either break the massive degeneracies by putting infinitesimal disorder which leads to the $S_{\text{quen}}$, or otherwise investigate quantities such as $S_{\text{typ}}$ and $S_{\text{ann}}$ which includes physics of the degenerate manifold. We uncover that such measures have contributions from both fermionic correlations and geometrical aspects. While $S_{\text{quen}}$ generically follow area law, $S_{\text{typ}}$ and $S_{\text{ann}}$ are volume law in character. These quantities can, in turn, be estimated from properties of the clusters, which either cut the entanglement bi-partition or contribute to the zero-mode degeneracies. Interestingly, the entanglement measure of the largest cluster can capture even the classical percolation threshold in two dimensions.
While in this work, we do not propose any experimental setups to measure such entanglement signatures, finding realistic proposals [79, 80, 81, 16] in this direction would be interesting to pursue. Similarly, we have focused in this study exclusively on bipartite entanglement entropy, given its relevance to quantum condensed matter systems. In quantum information, various other measures have been pursued to investigate phases and phase transitions [82, 83]. A comprehensive investigation of these, in regard to percolation disorder, is another prospective study. Finally, investigating this physics in both symmetry-protected topological systems and topologically ordered systems will be an exciting future direction.
Acknowledgement: We acknowledge fruitful discussions with Diptarka Das, Apratim Kaviraj, Saikat Ghosh, G. Sreejith, Saikat Mondal, and Harish Adsule. S.P. acknowledges funding from IIT Kanpur Institute Fellowship. AA acknowledges support from IITK Initiation Grant (IITK/PHY/2022010). Numerical calculations were performed on the workstation Wigner at IITK.
Supplemental Material to “Entanglement signatures of a percolating quantum system”
I Entanglement between two free spins
To understand the notion of typical and annealed entanglement in contrast with standard bipartite entanglement entropy, let us consider two free spins $A$ and $B$ as shown in Fig. S5(a). A general choice of the valid many-body wavefunction of the system will be,
$|\psi\rangle=a|\downarrow_{A}\downarrow_{B}\rangle+b|\downarrow_{A}\uparrow_{B}\rangle+c|\uparrow_{A}\downarrow_{B}\rangle+d|\uparrow_{A}\uparrow_{B}\rangle,$ (S1)
since the spins are free. Given the wavefunction coefficients (here $a,b,c$ and $d$) are complex random variables chosen from a Gaussian distribution with zero mean and unit variance, the random state $|\psi\rangle$ can be thought of as Haar uniform pure state [50, 51]. So, in a general form, any random pure state of our free spin system can be written as,
$|\psi\rangle=\sum_{i,j}\phi_{i,j}|\alpha_{A}^{i}\rangle\otimes|\alpha_{B}^{j}\rangle,$ (S2)
with $\alpha^{i},\alpha^{j}\in\{\uparrow,\downarrow\}$ is the basis of the Hilbert space and wavefunction coefficients $\phi_{i,j}$s (i.e. $a,b,c,d$) are the elements of a Gaussian random matrix $\phi$, satisfying the wavefunction normalisation constraint $\sum_{i,j}|\phi_{i,j}|^{2}=|a|^{2}+|b|^{2}+|c|^{2}+|d|^{2}=1$. Now, the partial density matrix for the spin A will be,
$\hat{\rho}_{A}$ $=$ $\text{Tr}_{B}(\hat{\rho})=\text{Tr}_{B}\big{(}\langle\psi|\psi\rangle\big{)}$ (S3)
$=$ $\sum_{m}{\langle\alpha_{B}^{m}|}\bigg{(}\sum_{i,j}\sum_{k,l}\phi_{i,j}\phi_{k,l}^{*}|\alpha_{A}^{i}\rangle{\langle\alpha_{A}^{k}|}\otimes|\alpha_{B}^{j}\rangle{\langle\alpha_{B}^{l}|}\bigg{)}|\alpha_{B}^{m}\rangle$
$=$ $\sum_{i,k}\sum_{m}\phi_{i,m}\phi^{*}_{k,m}|\alpha_{A}^{i}\rangle{\langle\alpha_{A}^{k}|}$
$=$ $\sum_{i,k}\chi_{i,k}|\alpha_{A}^{i}\rangle{\langle\alpha_{A}^{k}|},$
where $\chi_{i,k}=\sum_{m}\phi_{i,m}\phi^{*}_{k,m}$ is the matrix element of $\phi\phi^{\dagger}$, consequently $\hat{\rho}_{A}=\phi\phi^{\dagger}$ in the basis of $\alpha_{A}$. From this, it is straightforward to see in our system,
$\hat{\rho}_{A}=\begin{pmatrix}|a|^{2}+|b|^{2}&ac^{*}+bd^{*}\\ ca^{*}+db^{*}&|c|^{2}+|d|^{2}\end{pmatrix}.$ (S4)
Now using $\hat{\rho}_{A}$ one can numerically calculate bipartite entanglement entropy $S_{A}(\hat{\rho}_{A})=\text{Tr}\big{(}{\hat{\rho}_{A}\log\hat{\rho}_{A}}\big{)}$ and average it over numerous Gaussian ensembles of wavefunction coefficients {$a,b,c,d$} to find typical entanglement measure $S_{\text{typ}}=\langle S_{A}(\hat{\rho}_{A})\rangle\simeq 0.33$. We also define annealed entanglement as $S_{\text{ann}}=S_{A}(\langle\hat{\rho}_{A}\rangle)=\text{Tr}\big{(}{\langle\hat{\rho}_{A}\rangle\log\langle\hat{\rho}_{A}\rangle}\big{)}$. To calculate this, one first takes ensemble averaged (this average is also taken over wavefunction coefficients as described earlier) partial density matrix $\langle\hat{\rho}_{A}\rangle$ which turns out to be $\frac{1}{2}I_{2}$ here, and then find entanglement entropy to get $S_{\text{ann}}=\log 2$.
Figure S5: (a) Two free spin $A$ and $B$. (b) A four-site toy model with only two sites ($1$ and $4$) are connected via fermionic hopping of strength $t$. The bipartition used for entanglement calculation is shown in red-dashed line. (c) Different contributions in $S_{\text{typ}}$ for the four-site toy example. Contribution from fermionic correlation $S_{\text{corr}}$ is always $\log 2$. For complex random coefficients, the geometric contribution $S_{\text{geom}}$ converges to $S_{\text{geom}}(\mathbb{C})=\frac{1}{2}$ with increasing $N_{R}$ (number of random pure state ensembles), and for real choices it converges to $S_{\text{geom}}(\mathbb{R})=(2\log 2-1)$.
II Four site fermionic toy example
In order to illustrate how typicality in entanglement measure appears in a fermionic model, consider the system where two spinless fermions hops (with hopping strength $t=-1$) on a four-site structure as shown in Fig. S5(b). While the single-particle spectrum is straight-forwardly given by {±$\{\pm 1,0,0\}$, the many-body ground state is not unique but two-fold degenerate. This can be simply seen since, to minimize the energy, while one fermion resides in a delocalized state between sites 1 and 4, another fermion can either be on either site 2 or 3. The two valid ground-state wavefunctions of the system of the system are
$|\Psi_{1}\rangle$ $=$ $\frac{1}{\sqrt{2}}(c^{\dagger}_{1}+c^{\dagger}_{4})c^{\dagger}_{2}|\Omega\rangle,$ (S5)
$|\Psi_{2}\rangle$ $=$ $\frac{1}{\sqrt{2}}(c^{\dagger}_{1}+c^{\dagger}_{4})c^{\dagger}_{3}|\Omega\rangle,$ (S6)
where $|\Omega\rangle$ is the fermionic vacuum and the entanglement entropy across the partition is clearly $\log 2$ given the delocalized fermion. However, any general wavefunction of the form
$|\Psi\rangle=\phi_{1}|\Psi_{1}\rangle+\phi_{2}|\Psi_{2}\rangle,$ (S7)
is also a valid wavefunction of the system. Given the coefficients are randomly drawn from a Gaussian distribution and follow wavefunction normalisation constraint, they form a random pure state similar to what we discussed in I. This leads to similar considerations of typical entanglement $S_{\text{typ}}$ as described in the main text. However, now it has components arising from the fermionic correlations within the connected cluster across the partition (the delocalized state between site 1 and 4) and contributions due to a random choice of coefficients $\phi_{1}$ and $\phi_{2}$. In general,
$S_{\text{typ}}=S_{\text{corr}}+S_{\text{geom}},$ (S8)
where $S_{\text{corr}}$ captures the fermionic correlations in connected regions and $S_{\text{geom}}$ arises out of typicality due to degeneracy, often arising out of disconnected sites. Therefore, the latter has information about the geometrical properties of the network. In the present example, while $S_{\text{corr}}=\log 2$, $S_{\text{geom}}$ depends on the choices of the wavefunction coefficients ($\phi_{1}$ and $\phi_{2}$) used in the formation of random pure states, leading to the symmetry classification of the typical entanglement [55]. In class A [58], the coefficients are complex, and the analytical result shows $S_{\text{geom}}(\mathbb{C})=\frac{1}{2}$, while in class AI [58] where the coefficients are real, it is evaluated to be $S_{\text{geom}}(\mathbb{R})=(2\log 2-1)$. In Fig. S5(c), we show that our numerical results converge towards analytical results with the increasing number of pure ensembles ($N_{R}$) used in the calculation of $S_{\text{typ}}$.
Figure S6: $S_{\text{typ}}$ as a function of filling $f$, for the zero-dimensional fermionic problem. Numerical data is calculated for a system of $N=200$ sites with equal bipartition and averaged over $10^{4}$ ensembles of random pure states.
III Zero-dimensional fermionic problem
An all-to-all connected fermionic hopping network hosts $N_{0}=N-1$ number of zero energy states for a $N$ site system. So, a typical many-body wavefunction of the system will be a random pure state made out of these degenerate zero energy states. Here, we choose the wavefunction coefficients to be real, restricting it to class AI [55]. Now, $S_{\text{typ}}$ at half filling $f=\frac{1}{2}$, and for equal bipartition, typical entanglement contain a small fermionic correlation part and a huge geometric contribution because of the massive degeneracy in the spectrum. While the geometric part follows the symmetry classification result [55], the fermionic part is captured in $S_{\text{quen}}$, evaluated using the degeneracy breaking protocol as discussed in the main text. Now for any general filling $f$ ($10<f<$ 1) with equal bipartition, the available phase space volume is $f(1-f)$; consequently, the effective degenerate Hilbert space dimension becomes $\sim 2^{af(1-f)N_{0}}$, where $a$ is a dimensionless constant. Then, using Page’s result [48], we empirically find the typical entanglement in the leading order to be,
$S_{\text{typ}}(f)=S_{\text{quen}}(f)+4f\big{(}1-f\big{)}\big{(}N_{0}-1\big{)}\big{(}\log 2-\frac{1}{2}\big{)},$ (S9)
so that it remains consistent with the result at half filling ($S_{\text{typ}}=S_{\text{quen}}+\big{(}N_{0}-1\big{)}\big{(}\log 2-\frac{1}{2}\big{)}$). In Fig. S6, we show the numerical values as a function of filling ($f$), which match well with our empirical formula.
IV Lattice animals, their entanglement and number of zeromodes
In the case of square lattice percolation, the quenched entanglement $S_{\text{quen}}$ can be captured using the lattice animals and their contributions to entanglement given a bipartition. For a lattice animal of bond size $k$ (i.e. number of bonds in that lattice animal is $k$) and perimeter $t$, we calculate the bipartite entanglement entropy $S_{\text{animal}}^{b}$ with the partition dissecting the animal in a certain way. Now if $P_{\text{animal}}^{b}$ is the probability that the different animals of the same {$k,t$} appear on that given partition and give rise to the same $S_{\text{animal}}^{b}$, then for a square lattice of linear size $L$,
$S_{\text{quen}}\approx\sum_{\text{animal},b}P_{\text{animal}}^{b}S_{\text{animal}}^{b}L.$ (S10)
The factor of $L$ comes because there are $L$ number of sites on the partition where the animal can appear.
Similarly, we calculate the number of zero energy modes in the percolating square lattice using lattice animal enumeration. The mean number of bond clusters is given by $g(k,t)p^{k}(1-p)^{t}$ and $g(k,t)$ is the number of lattice animals of size $\{k,t\}$ [77]. Then the number of zeromodes for $L$ sized system is,
$N_{0}=\sum_{k,t}n_{0}(k,t)g(k,t)p^{k}(1-p)^{t}L^{2},$ (S11)
where $n_{0}(k,t)$ is the number of zeromodes in a cluster of size $\{k,t\}$. For small values of $k$ (up to $4$) and $t$ (up to $12$), all these data are tabulated in TABLE 1. Now, using eqn. (S10) and eqn. (S11), one can calculate quenched entanglement and the number of zeromodes in the percolated square lattices. Since these small-sized clusters are dominant in small values of percolation probability $p$, our result from TABLE 1 matches the disorder averaged numerical results ($\tilde{S}_{\text{quen}}$ and $\tilde{N}_{0}$) at small $p$ values (see main text).
Table 1: Small-sized lattice animals (up to bond-size $k=4$ and perimeter $t=12$) and their contribution in quenched entanglement $S_{\text{quen}}$ and number of zeromodes $N_{0}$ for percolated square lattice. $S_{\text{animal}}^{b}$ is the entanglement of a lattice animal given the bipartition cuts the animal in a certain way, and $P_{\text{animal}}^{b}$ is the probability of that animal appears on that bipartition. The prefactors in $P_{\text{animal}}^{b}$ are decided by the number of animal structures with the same partition, giving the same $S_{\text{animal}}^{b}$. For a given {$k,t$}, $g(k,t)$ is the number of lattice animals, and $n_{0}(k,t)$ is the number of zero energy states they can host.
Animals $S_{\textbf{animal}}^{b}$ $P_{\textbf{animal}}^{b}$ $k,t$ $\bm{g(k,t)}$ $n_{0}(k,t)$
00 $q^{4}$ $0,4$ $1$ $1$
$\log 2$ $pq^{6}$ $1,6$ $2$ 00
$0.56234$ $6p^{2}q^{8}$ $2,8$ $6$ $1$
$\log 2$ $14p^{3}q^{10}$
$0.41328$ $p^{3}(7q^{10}+2q^{9})$ $3,10$ $14$ 00
$0.45056$ $6p^{3}q^{10}$ $3,10$ $4$ $1$
$1.17903$ $2p^{3}q^{9}$ $3,9$ $4$ 00
$0.54387$ $p^{4}(34q^{12}+8q^{11})$
$0.56234$ $p^{4}(20q^{12}+8q^{11})$
$\log 2$ $6p^{4}q^{12}$
$1.14317$ $2p^{4}q^{12}$ $4,12$ $54$ $1$
$1.17349$ $8p^{4}q^{11}$
$0.95494$ $8p^{4}q^{11}$
$0.32346$ $p^{4}(10q^{12}+8q^{11})$
$0.63651$ $p^{4}(30q^{12}+8q^{11})$ $4,11$ $32$ $1$
$0.37677$ $2p^{4}q^{12}$ $4,12$ $1$ $3$
$0.83299$ $p^{4}q^{8}$ $4,8$ $1$ $2$